OpenStudy (anonymous):

Which algebraic expression would not be equivalent if the number and the variable were reversed? A.n + 8 B.m – 5 C.k x 8 D.4 + x

OpenStudy (anonymous):

D?

OpenStudy (austinl):

Well, I would say B. \(n+8\Rightarrow2+8=8+2\\ m-5\Rightarrow2-5\ne5-2\\ k\times8\Rightarrow2\times8=8\times2\\ 4+x\Rightarrow4+2=2+4\)

OpenStudy (anonymous):

that makes sense thanks :)

OpenStudy (austinl):

You are very welcome!